# Advancing Semantic Textual Similarity Modeling: A Regression Framework with Translated ReLU and Smooth K2 Loss

### v1

@@ Abstract
Since the introduction of BERT and RoBERTa, research on Semantic Textual Similarity (STS) has made groundbreaking progress. Particularly, the adoption of contrastive learning has substantially elevated state-of-the-art performance across various STS benchmarks. However, contrastive learning categorizes text pairs as either semantically similar or dissimilar, failing to leverage fine-grained annotated information and necessitating large batch sizes to prevent model collapse. These constraints pose challenges for researchers engaged in STS tasks that require nuanced similarity levels or those with limited computational resources, compelling them to explore alternatives like Sentence-BERT. Nonetheless, Sentence-BERT tackles STS tasks from a classification perspective, overlooking the progressive nature of semantic relationships, which results in suboptimal performance. To bridge this gap, this paper presents an innovative regression framework and proposes two simple yet effective loss functions: Translated ReLU and Smooth K2 Loss. Experimental analyses demonstrate that our method achieves convincing performance across seven established STS benchmarks, especially when supplemented with task-specific training data. 111Our code and checkpoints are available at https://github.com/ZBWpro/STS-Regression.
@@ 1 Introduction
Semantic Textual Similarity (STS) constitutes a fundamental task in natural language processing, wielding significant influence across a multitude of applications including text clustering, information retrieval, and recommendation systems. Despite the remarkable precision achieved by interactive architectures within these tasks, their inability to support offline computation limits their viability in large-scale text analysis scenarios. In response to this, the seminal work of Sentence-BERT (Reimers and Gurevych, 2019) introduces a dual-tower architecture to encode the sentences within a pair separately, thereby facilitating the derivation of independent embeddings. This approach showcases superior efficacy and has rapidly gained widespread acceptance, now serving as a cornerstone for various downstream tasks. Consequently, further improvements to Sentence-BERT hold high research and practical value.
@@
Nevertheless, the advent of contrastive learning methods, exemplified by SimCSE (Gao et al., 2021), has demonstrated more pronounced enhancements across renowned English STS benchmarks like STS12-16 (Agirre et al., 2012, 2013, 2014, 2015, 2016), STS-B (Cer et al., 2017), and SICK-R (Marelli et al., 2014). This has shifted the research focus in recent years towards integrating contrastive learning techniques with pre-trained language models (PLMs) like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019). An intuitive comparison is that using the NLI dataset (Bowman et al., 2015; Williams et al., 2018) as a training corpus, SimCSE-RoBERTa attains an average Spearman’s correlation score of 82.52 across these STS tasks, hugely surpassing the 74.21 achieved by Sentence-RoBERTa.
@@
Such discernible performance disparity has inadvertently overshadowed the advantages of Sentence-BERT, especially in terms of data utilization efficiency and computational resource demands. Contrastive learning, by its self-supervised nature, predominantly recognizes text pairs as either similar or dissimilar. This binary categorization restricts contrastive learning methods to using triple form data composed of an anchor sentence, a positive instance, and a hard negative instance for training in supervised settings (Gao et al., 2021). Many practical scenarios, however, tend to provide more finely-grained labeled data (e.g., highly relevant, moderately relevant, relevant, not relevant) (Liu et al., 2023), where contrastive learning approaches can usually only exploit text pairs whose similarity indicators are at the endpoints.
@@
Additionally, since contrastive learning enhances model discriminability by treating other samples within the same batch as negative instances, it requires large batch sizes, thereby consuming substantial computational resources. For example, SimCSE’s supervised learning settings include a batch size of 512 and 3 epochs. To accommodate this configuration on consumer-grade GPUs, SimCSE constrains the maximum input length to 32 (Gao et al., 2021). In contrast, Sentence-BERT and our proposed methodology necessitate a mere batch size of 16 and 1 epoch to reach convergence. Additionally, our default maximum input length is 256, significantly longer than SimCSE’s.
@@
The aforementioned drawbacks highlight the difficulty in completely replacing Sentence-BERT with contrastive learning methods. Hence, some cutting-edge works (Zhang et al., 2023b) continue to employ Sentence-BERT for sentence embedding derivation. Nonetheless, given that STS tasks typically categorize text pairs by degrees of semantic similarity, and Sentence-BERT approaches these tasks from a classification standpoint, neglecting the progressive relationships between categories, there exists a clear opportunity for improvement. As a illustration, consider an STS task with five categories, labeled consecutively from 1 to 5. Traditional classification strategies would yield identical loss for a sample scored at 2, irrespective of its prediction as 3 or 4, an approach evidently suboptimal.
@@
To rectify such deficiency, this paper proposes a novel framework that converts multi-category STS tasks into regression problems, thus effectively capturing the progressive relationships between categories. For a given dataset, we first map its original labels to a sequential array of integers, ensuring that samples with higher similarity scores are assigned correspondingly greater integers. Then, we set the number of nodes in the output layer to one, thereby enabling the model to produce a continuous prediction value. Finally, the model parameters are updated according to the difference between predicted and actual scores.
@@
Distinct from standard regression problems, the ground truth within our transformed multi-category STS tasks manifest as a series of discrete points along the numerical axis. Therefore, instead of precisely matching the target points, the floating-point predictive values just need to be sufficiently close to get correctly classified. To accommodate this process, we introduce a zero-gradient buffer zone to widely utilized L1 Loss and MSE Loss, unveiling two innovative loss functions: Translated ReLU and Smooth K2 Loss.
@@
Comprehensive evaluations across seven STS benchmarks substantiate that our regression framework surpasses traditional classification strategies in handling multi-category STS tasks. Additionally, we find that further updating our model’s checkpoint with the STS-B and SICK-R training sets allows our method to achieve superior Spearman correlation relative to contrastive learning methods, reaching state-of-the-art performance. These findings reinforce the effectiveness of our proposed solution and the importance of utilizing task-specific data, an aspect often neglected in contrastive learning paradigms.
@@
We propose two innovated loss functions, Translated ReLU and Smooth K2 Loss, specifically tailored to address classification problems involving progressive relationships between categories.
Through empirical evidence, we demonstrate that, when combined with task-specific data, our Siamese network approach can attain better results than contrastive learning schemes.
@@ 2 Related Work
In this chapter, we primarily review two types of STS task solutions directly related to our work:
Siamese Neural Network Architectures: These approaches (Reimers and Gurevych, 2019; Conneau et al., 2017; Thakur et al., 2021), proposed relatively earlier in the field, have been widely applied across various domains owing to their effectiveness on annotated data. Although their performance on the seven STS benchmarks (STS 12-16, STS-B, SICK-R) is generally inferior to contemporary contrastive learning methods, this discrepancy largely arises from the absence of task-specific training data. Thus, models have the flexibility to opt for alternative sources, such as wiki datasets (Gao et al., 2021) or NLI datasets (Bowman et al., 2015; Williams et al., 2018), which adapt readily to triplet format. Given our goal of tackling multi-category STS tasks, our model architecture remains rooted in the Siamese network. However, in contrast to preceding efforts, we introduce an innovative regression framework designed to explicitly capture the progressive relationships between categories.
Contrastive Learning Methods: Contrastive learning has become the dominant paradigm for addressing STS tasks, characterized by a vast amount of research (Jiang et al., 2022; Zhang et al., 2023a). However, contrastive learning loss functions, epitomized by InfoNCE Loss (Oord et al., 2018), concentrate solely on the binary semantic categorization of texts and cannot directly utilize fine-grained labeled corpus. Furthermore, the necessity for large batch sizes to ensure negative sample diversity and prevent model collapse imposes considerable computational demands. For instance, supervised SimCSE’s training requires of GPU memory (Jiang et al., 2023), whereas our proposed method, even with a maximum sequence length eight times that of SimCSE, demands merely .
@@ 3 Methodology
This chapter delineates our methodological framework, beginning with a detailed exposition of the designed network architecture and its operational workflow in Section 3.1. Then, in Sections 3.2 and 3.3, we present the two novel loss functions proposed in this study.
@@ 3.1 Network Architecture
As illustrated in Figure 1, we utilize a Siamese neural network with shared parameters for encoding input sentences via BERT to obtain corresponding word embedding matrices. Subsequently, sentence embeddings, denoted as and for paired sentences and , are derived through average pooling. These embeddings, both vectors of the hidden dimension, are then concatenated alongside their element-wise difference and passed through a fully connected layer with parameters sized at to produce the model’s predicted continuous similarity score.
@@
Our methodology diverges from the original dual-tower structures employed by Sentence-BERT and InferSent (Conneau et al., 2017) in three critical aspects:
@@
1. We model STS tasks, characterized by a progressive relationship between categories, as regression problems. This is achieved by mapping labels from the original dataset to a sequence of incrementing integers reflective of their similarity relations, thus conveying to the model that categories are not independent but progressively related.
@@
2. Building on this, we streamline the output node count in our final fully connected layer to one, enabling the model to directly yield a similarity score rather than a categorical probability distribution. Through this adjustment, for STS tasks containing categories, we effectively reduce the parameter size of the output layer from to . In light of the expanding dimensions of hidden layers in contemporary PLMs, this optimization can save substantial computational resources.
@@
3. Contrasting with InferSent and Sentence-BERT’s classification-based approach, which assigns target classes for sentence pairs based on the highest probability, our regression framework categorizes based on the closeness between the predicted and actual values.
@@
To better understand this process, consider an STS task with four categories, labeled as “very relevant,” “moderately relevant,” “slightly relevant,” and “not relevant.” After clarifying the progressive relationship between these categories, we would map them to four consecutive integers 0, 1, 2, 3, respectively, ranging from “not relevant” to “very relevant.” This mapping strategy is highly flexible, allowing for task-specific adjustments in numerical nodes and intervals. Subsequently, we encode the paired sentences separately and calculate their semantic similarity, resulting in a floating-point prediction value. By rounding this value, it can be converted into a discrete label. For instance, a prediction of 2.875 for a sample pair would be classified as “very relevant,” as it approximates closely to the boundary point 3. Similarly, if a sample’s prediction is 1.333, it would be approximated to 1 and thus classified as “slightly relevant" because 1.333 is closer to 1 among the four boundary points 0, 1, 2, 3.
@@
Extending from the above examples, it can be seen that if we map the original labels to natural numbers spaced by , as long as the difference between the model’s prediction and the ground truth is less than , the sample will be correctly classified. However, conventional regression loss functions, represented by L1 Loss and MSE Loss, always enforce the difference between the model’s prediction and the true value to be zero—a requirement that is unnecessary for our scenario. Thus, we introduce a zero-gradient buffer zone into both functions, resulting in the creation of Translated ReLU and Smooth K2 Loss.
@@ 3.2 Translated ReLU
We first present Translated ReLU, mathematically formulated in Equation 1. Herein, represents the interval between mapped category labels, with for a sequence of consecutive natural numbers.
As previously discussed, when the difference between the model’s predicted value and the ground truth is less than , it signifies a correct classification of the sample. Traditional regression loss functions, however, mandate absolute congruence between predictions and true values, applying a penalty for any deviation. This stringent requirement to some extent diverts the model’s focus from difficult samples that have not yet been correctly classified and ignores the inherent variability within classes.
@@ 3.3 Smooth K2 Loss
Echoing the structure of Translated ReLU, Smooth K2 Loss also incorporates a zero-gradient buffer zone, but exhibits a quadratic function for , as illustrated on the right side of Figure 2. Given the differential mathematical underpinnings of these two loss functions, Smooth K2 Loss is recommended for scenarios with high-quality data and strong credibility. In contrast, when dealing with datasets that contain considerable noise, Translated ReLU may be a more suitable choice.
@@ 4 Experiment
This chapter provides empirical validation of our regression framework and two innovative loss functions: Translated ReLU and Smooth K2 Loss. We commence by comparing the performance across different modeling strategies for multi-category STS tasks and various loss functions (Section 4.1). Subsequently, we demonstrate that, when supplemented with task-specific training data, our Siamese neural network architecture outperforms prevailing contrastive learning methods (Section 4.2). Following this, we examine the influence of varying hyperparameter settings on model performance (Section 4.3). Finally, we present ablation studies of our proposed methodology (Section 4.4).
4.1 STS Performance without Task-specific Training Data
Our experimental setup closely mirrors that of Sentence-BERT, leveraging fine-tuning on BERT or RoBERTa with a composite corpus derived from SNLI and MNLI datasets. These NLI datasets categorize sentence pairs into three distinct classes: contradiction, neutral, and entailment. Sentence-BERT maps these to 0, 2, 1, respectively, and employs a classification strategy for training (Reimers and Gurevych, 2019). In contrast, our method sequentially maps contradiction, neutral, and entailment to 0, 1 and 2. This mapping reflects the natural order of semantic similarity, from least to most similar, thereby enabling our regression framework to more effectively capture the progressive relationships between categories.
For computational efficiency, we uniformly set the batch size to 16 and limit training to a single epoch, with model checkpoints preserved based on performance metrics on the STS-B development set. The specific hyperparameter settings for Translated ReLU and Smooth K2 Loss are cataloged in Table 1. During the evaluation phase, we assess the model’s average Spearman correlation across seven STS tasks via the SentEval (Conneau and Kiela, 2018) toolkit. The results of the aforementioned experiments are summarized in Table 2, from which we distill insights along three pivotal aspects:
| PLM | Loss | ||
| BERT | Translated ReLU | 2.5 | 0.25 |
| BERT | Smooth K2 Loss | 2 | 0.25 |
| RoBERTa | Translated ReLU | 1 | 0.25 |
| RoBERTa | Smooth K2 Loss | 3 | 0.25 |
1. Classification Strategy vs. Regression Strategy: Our regression framework, particularly when utilizing Smooth K2 Loss, yields an average Spearman correlation of 76.03 for BERT and 76.04 for RoBERTa. These figures significantly outstrip those attained through Sentence-BERT and the classification method with Cross-Entropy Loss, highlighting the regression-based modeling’s superiority in both reducing the output layer’s parameter size and enhancing semantic discrimination for multi-category STS tasks.
2. Efficacy of the Zero-Gradient Buffer Zone: The adoption of Translated ReLU improves performance for both BERT and RoBERTa beyond what is achieved with L1 Loss. Likewise, employing Smooth K2 Loss surpasses MSE Loss on both PLMs. These comparisons underline the benefit of integrating a zero-gradient buffer zone in balancing model’s focus across diverse samples within regression-modeled multi-category classification tasks.
3. Adaptive Gradients Aligned with Prediction Errors: Models trained with Smooth K2 Loss outshine those utilizing Translated ReLU, and models employing MSE Loss exceed those with L1 Loss. This evidences the advantages of dispensing differentiated gradients in line with prediction-ground truth deviations, especially when leveraging high-quality datasets like NLI.
Collectively, these findings substantiate the merit of adopting a regression framework for multi-category STS tasks and enhancing traditional regression loss functions with a zero-gradient buffer zone to optimize model performance.

### v2

@@ Abstract
Since the introduction of BERT and RoBERTa, research on Semantic Textual Similarity (STS) has made groundbreaking progress. Particularly, the adoption of contrastive learning has substantially elevated state-of-the-art performance across various STS benchmarks. However, contrastive learning categorizes text pairs as either semantically similar or dissimilar, failing to leverage fine-grained annotated information and necessitating large batch sizes to prevent model collapse. These constraints pose challenges for researchers engaged in STS tasks that involve nuanced similarity levels or those with limited computational resources, compelling them to explore alternatives like Sentence-BERT. Despite its efficiency, Sentence-BERT tackles STS tasks from a classification perspective, overlooking the progressive nature of semantic relationships, which results in suboptimal performance. To bridge this gap, this paper presents an innovative regression framework and proposes two simple yet effective loss functions: Translated ReLU and Smooth K2 Loss. Experimental results demonstrate that our method achieves convincing performance across seven established STS benchmarks and offers the potential for further optimization of contrastive learning pre-trained models. 111Our code and checkpoints are available at https://github.com/ZBWpro/STS-Regression.
@@ 1 Introduction
Semantic Textual Similarity (STS) constitutes a fundamental task in natural language processing, wielding significant influence across a multitude of applications, including text clustering, information retrieval, and recommendation systems. Despite the remarkable precision obtained by interactive architectures within these tasks, their inability to support offline computation limits their viability in large-scale text analysis scenarios. In response, the seminal work of Sentence-BERT (Reimers and Gurevych, 2019) introduces a dual-tower architecture to encode the sentences within a pair separately, thereby facilitating the derivation of independent embeddings. This approach showcases superior efficacy and has rapidly gained widespread acceptance, now serving as a cornerstone for various downstream tasks. Consequently, further improvements to Sentence-BERT hold significant research interest and practical value.
@@
Nevertheless, the advent of contrastive learning methods, exemplified by SimCSE (Gao et al., 2021), has led to more pronounced enhancements on renowned English STS benchmarks, such as STS12-16 (Agirre et al., 2012, 2013, 2014, 2015, 2016), STS-B (Cer et al., 2017), and SICK-R (Marelli et al., 2014). This has shifted the research focus in recent years towards integrating contrastive learning techniques with pre-trained language models (PLMs) like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019). An intuitive comparison is that, when employing the NLI dataset (Bowman et al., 2015; Williams et al., 2018) as a training corpus, SimCSE-RoBERTa attains an average Spearman’s correlation score of 82.52 across these STS tasks, hugely surpassing the 74.21 achieved by Sentence-RoBERTa.
@@
Such discernible performance disparity has inadvertently overshadowed the advantages of Sentence-BERT, especially in terms of data utilization efficiency and computational resource demands. Contrastive learning, by its self-supervised nature, predominantly recognizes text pairs as either similar or dissimilar. This binary categorization restricts contrastive learning methods to training on triplet-form data composed of an anchor sentence, a positive instance, and a hard negative instance in supervised settings (Gao et al., 2021). Many practical scenarios, however, tend to provide more finely grained labeled data (e.g., highly relevant, moderately relevant, relevant, and irrelevant) (Liu et al., 2023), where contrastive learning approaches can usually only exploit text pairs whose similarity indicators are at the endpoints.
@@
Furthermore, since contrastive learning enhances model discriminability by treating other samples within the same batch as negative instances, it requires large batch sizes, thereby consuming substantial computational resources. For example, SimCSE’s supervised learning settings include a batch size of 512 and 3 epochs. To accommodate this configuration on consumer-grade GPUs, SimCSE limits the maximum input length to 32 tokens (Gao et al., 2021). In contrast, Sentence-BERT and our proposed methodology necessitate a mere batch size of 16 and 1 epoch to reach convergence. Additionally, our default maximum input length is 256, significantly longer than SimCSE’s.
@@
The aforementioned drawbacks highlight the difficulty in completely replacing Sentence-BERT with contrastive learning methods. Hence, some cutting-edge works (Zhang et al., 2023) continue to rely on Sentence-BERT for sentence embedding derivation. Nonetheless, given that STS tasks typically categorize text pairs by degrees of semantic similarity, and Sentence-BERT approaches these tasks from a classification standpoint, neglecting the progressive relationships between categories, there exists a clear opportunity for improvement. As an illustration, consider an STS task with five categories, labeled consecutively from 1 to 5. Traditional classification strategies would yield identical loss for a sample scored at 2, irrespective of its prediction as 3 or 4, an approach evidently suboptimal.
@@
To rectify such deficiency, this paper proposes a novel framework that converts multi-class STS tasks into regression problems, thus effectively capturing the progressive relationships between categories. For a given dataset, we first map its original labels to evenly spaced numerical values, ensuring that samples with higher similarity scores are assigned correspondingly greater values. Then, we set the number of nodes in the output layer to one, thereby enabling the model to produce a continuous prediction. Finally, the model parameters are updated according to the difference between predicted and actual scores.
@@
Distinct from standard regression tasks, the ground truth within our transformed multi-category STS tasks manifest as a series of discrete points along the numerical axis. Therefore, instead of requiring precise matches to the target values, the floating-point predictions just need to be sufficiently close to get correctly classified. To accommodate this characteristic, we introduce a zero-gradient buffer zone to widely utilized L1 Loss and MSE Loss, unveiling two innovative loss functions: Translated ReLU and Smooth K2 Loss.
@@
Comprehensive evaluations across seven STS benchmarks substantiate that our regression framework surpasses traditional classification strategies in handling multi-category STS tasks. Additionally, we find that our approach can further refine the performance of contrastive learning pre-trained models by utilizing filtered STS-B and SICK-R training sets. These findings highlight the effectiveness of our method and underscore the importance of harnessing task-specific data, an aspect often neglected in contrastive learning paradigms.
@@
We propose two novel loss functions, Translated ReLU and Smooth K2 Loss, specifically tailored to address classification problems involving progressive relationships between categories.
Through empirical evidence, we demonstrate that our strategy can be combined with leading contrastive learning pre-trained models, leveraging fine-grained annotated data to further improve their performance. This offers a new perspective for current research in STS and sentence embeddings.
@@ 2 Related Work
In this section, we primarily review three types of STS solutions that are directly relevant to our work:
Siamese Neural Network Architectures: These approaches (Reimers and Gurevych, 2019; Conneau et al., 2017; Thakur et al., 2021), proposed relatively earlier in the field, have been widely applied across various domains owing to their effectiveness on annotated corpus. Although their performance on the seven STS benchmarks (STS 12-16, STS-B, SICK-R) is generally inferior to contemporary contrastive learning methods, this disparity largely stems from the absence of task-specific training data. Thus, models have the flexibility to opt for alternative sources, such as Wikipedia (Gao et al., 2021) or NLI datasets (Bowman et al., 2015; Williams et al., 2018), which adapt readily to triplet format. Given our goal of tackling multi-category STS tasks, our model architecture remains rooted in the Siamese network. However, in contrast to preceding efforts, we introduce an innovative regression framework specifically designed to capture the progressive relationships between categories.
Contrastive Learning Fine-Tuning Methods: Contrastive learning is currently the mainstream paradigm for addressing STS tasks, with substantial research exploring its integration with the fine-tuning of PLMs (Jiang et al., 2022a; Zhang et al., 2024). However, contrastive learning loss functions, epitomized by InfoNCE Loss (Oord et al., 2018), concentrate exclusively on binary semantic categorization and are unable to fully utilize fine-grained labeled texts. Additionally, the necessity for large batch sizes to ensure negative sample diversity and prevent model collapse imposes significant computational demands. These two limitations are inherently difficult to overcome within contrastive learning itself, yet they are precisely the strengths of Sentence-BERT-style dual-tower models. Therefore, a primary objective of this paper is to investigate whether the performance of contrastive learning models can be further enhanced by incorporating traditional Siamese neural network architectures.
Contrastive Learning Pre-Trained Models: With the growing importance of embeddings in retrieval-augmented generation (Zhao et al., 2024) and other application scenarios, more companies and institutions are dedicating efforts to developing specialized text representation models. These approaches generally adopt multi-stage contrastive learning strategies for network pre-training (Wang et al., 2022; Li et al., 2023; Xiao et al., 2024). Additionally, compared to large-scale generative PLMs, lightweight discriminative models that capture bidirectional dependencies are often more preferred. In our experiments section, we employ two state-of-the-art contrastive learning pre-trained models, Jina Embeddings v2 (Günther et al., 2023) and Nomic Embed (Nussbaum et al., 2024). Both are BERT-based encoder architectures with a parameter size of 137 million.
@@ 3 Methodology
This section presents our methodological framework, beginning with a detailed exposition of the network architecture and its operational workflow in subsection 3.1. Then, in subsections 3.2 and 3.3, we introduce the two novel loss functions proposed in this study.
@@ 3.1 Network Architecture
As illustrated in Figure 1, we utilize a Siamese neural network with shared parameters for encoding input sentences via BERT to obtain corresponding word embedding matrices. Subsequently, sentence embeddings, denoted as and for paired sentences and , are derived through average pooling. These embeddings, both vectors of the hidden dimension, are then concatenated alongside their element-wise difference and passed through a fully connected layer with parameters sized at to produce the model’s predicted similarity score.
@@
Our method diverges from the original dual-tower structures employed by Sentence-BERT and InferSent (Conneau et al., 2017) in three critical aspects:
@@
1. We model STS tasks, characterized by a progressive relationship between categories, as regression problems. This is achieved by mapping labels from the original dataset to a sequence of incrementing numbers reflective of their similarity relations, thus conveying to the model that categories are not independent but progressively related.
@@
2. Building on this, we streamline the output node count in the final fully connected layer to one, thereby enabling the model to directly yield a similarity score rather than a categorical probability distribution. Through this adjustment, for STS tasks with categories, we effectively reduce the parameter size of the output layer from to . In light of the expanding hidden layer dimensions in modern PLMs, this optimization can save considerable computational resources.
@@
3. Unlike the classification-based approach of InferSent and Sentence-BERT, which assigns target classes for sentence pairs according to the highest probability, our regression framework categorizes based on the closeness between the predicted and actual values.
@@
To better understand this process, consider an STS task with four categories: “highly relevant,” “moderately relevant,” “slightly relevant,” and “irrelevant.” After clarifying the progressive relationship between these categories, we would map them to four consecutive numbers 0, 1, 2, 3, respectively, ranging from “irrelevant” to “highly relevant.” This mapping strategy is flexible, allowing for task-specific adjustments in both numerical nodes and intervals. Furthermore, the mapped nodes do not necessarily have to be integers. Subsequently, we encode the paired sentences and compute their semantic similarity, resulting in a floating-point prediction. By rounding this value, it can be converted into the corresponding label. For instance, a prediction of 2.875 for a sample pair would be classified as “highly relevant,” as it is closest to the boundary point of 3. Similarly, if a sample receives a predicted value of 1.333, it would be approximated to 1 and thus classified as “slightly relevant," because 1.333 is closer to 1 among the four boundary points 0, 1, 2, 3.
@@
Extending from the above examples, it can be seen that if the original labels are mapped to nodes spaced by , as long as the difference between the model’s prediction and the ground truth is less than , the sample will be correctly classified. Specifically, for consecutive natural numbers, is equal to 1. However, conventional regression loss functions, represented by L1 Loss and MSE Loss, always enforce the model to exactly match the true value, a requirement that is unnecessary for our task scenario. Thus, we introduce a zero-gradient buffer zone into both functions, unveiling two new loss functions: Translated ReLU and Smooth K2 Loss.
@@ 3.2 Translated ReLU
We first present Translated ReLU, mathematically formulated in Equation 1. Herein, represents the interval between mapped category labels. As previously discussed, when the difference between the model’s predicted value and the ground truth is less than , it signifies a correct classification of the sample. Traditional regression loss functions, however, mandate absolute congruence between predictions and true values, applying a penalty for any deviation. This stringent requirement to some extent diverts the model’s focus from difficult samples that have not yet been correctly classified and ignores the inherent variability within classes.
@@ 3.3 Smooth K2 Loss
Echoing the design of Translated ReLU, Smooth K2 Loss also incorporates a zero-gradient buffer zone, but exhibits a quadratic function for , as illustrated on the right side of Figure 2. Given the differential mathematical underpinnings of these two loss functions, Smooth K2 Loss is recommended for scenarios with high-quality data and strong credibility. In contrast, when dealing with datasets that contain considerable noise, Translated ReLU may be a more suitable choice.
@@ 4 Experiment
This section provides empirical validation of our regression framework and two innovative loss functions. We commence by comparing the performance of different modeling strategies for multi-category STS tasks and various loss functions (subsection 4.1). Next, we demonstrate that, when supplemented with fine-grained training data, our Siamese neural network can effectively enhance the performance of contrastive learning PLMs (subsection 4.2). Following this, we highlight the computational efficiency of our methodology (subsection 4.3) and explore the influence of varying hyperparameter settings on model performance (subsection 4.4). Finally, subsection 4.5 presents ablation studies on our network architecture.
4.1 STS Performance Based on Traditional Discriminative Pre-Trained Models